OpenStudy (anonymous):

The formula, d = vt + 0.5at2, can be used to find the distance (d) traveled by an object, where v represents the initial velocity of the object, t represents the time, in seconds, and a represents the acceleration, in meters per second squared. A bicycle traveling at a velocity of 4.5 meters per second accelerates down a hill at a constant rate of 0.4 meters per second squared for 12 seconds. What is the distance, in meters, that the bicycle travels?

OpenStudy (anonymous):

d= vt = .5 at ^2

OpenStudy (anonymous):

Simply substitute v = 4.5 m/sec t = 12 sec a = 0.4 m/sec^2 d = (4.5)(12) + 0.5(0.4)(12^2) d = 54 +28.8 d = 82.8 meters :)